OpenStudy (anonymous):

\in a right triangle find the elngth of the side not given b=2 c=[\sqrt{?}\]29

OpenStudy (anonymous):

if c is the hypothenuns, then the length of the missing side is 5.

OpenStudy (anonymous):

a^2+b^2=c^2 so if c^2=29 29-4=a^2=25 a=5

OpenStudy (katielong):

 example: